if y times 1 minus x=z than what is x?
Found 2 solutions by solver91311, RAY100:
Answer by solver91311(24713)





Multiply by :



Add -1 to both sides:



Multiply by -1, then apply the LCD of on the right:



Remember to exclude zero as a possible value for , so:



That's as far as you can go without some additional information.

John


Answer by RAY100(1637)

y * (1-x) = z
.
y-xy =z
.
-xy = z -y
.
x = (z-y) / ( -y) = -z/y +1 =1-z/y
.
check by letting x=4, y=2, then z=-6,,,,ok for both original and answer